OpenStudy (bossimbacon):

Which expressions would complete this equation so that the equation has no solution? 6(4x – 3) – 10x = Choose exactly two answers that are correct. A. 14x – 15 B. 2(7x – 10) C. 15x – 14 D. 2(7x – 9) @TheSmartOne @immanuelv

OpenStudy (immanuelv):

there are two answers?

TheSmartOne (thesmartone):

Distribute

OpenStudy (bossimbacon):

yup only 2 answers

OpenStudy (studygurl14):

okay, so ya know how i gave you wrong answers before...well that's cause i misread the problem...now i understand it, so yea, i can help you better...lol

OpenStudy (bossimbacon):

i seee :D

OpenStudy (studygurl14):

But first, do what @TheSmartOne said...

TheSmartOne (thesmartone):

24x-18-10x=??

OpenStudy (studygurl14):

^^^

OpenStudy (immanuelv):

*hint* the solutions with same slope wont be the answer :D

OpenStudy (studygurl14):

^ that's a good hint

OpenStudy (studygurl14):

but not always necessarily true

OpenStudy (bossimbacon):

YOU GUYS ARE KILLING ME ITS HALLOWEEN IM DOING HW BE MY CANDY AND ANSWER THIS PLEASE

OpenStudy (studygurl14):

or is it...hold on

OpenStudy (immanuelv):

unless they have the same y intercept.. but not here xD

OpenStudy (immanuelv):

im below the food chain @Bossimbacon i cant disobey Mathletes xD

OpenStudy (studygurl14):

no...actually for this one, the slopes can be the same...you are finding ones that will produce no solutions, not infinte

OpenStudy (studygurl14):

lol @immanuelv

OpenStudy (studygurl14):

examples: 4x + 1 = 4x +2 No solution

OpenStudy (studygurl14):

4x +1 = 4x +1 inifnite solutions

OpenStudy (immanuelv):

haha well its the ones with same slope.b for the answer :d

OpenStudy (immanuelv):

i have put the answers in my answer as a treat for ya :D

OpenStudy (bossimbacon):

THANKK YOU SO MUCH @immanuelv so b and d'

OpenStudy (studygurl14):

A is one of the answers becaue \[14x - 18 \neq 14x - 15\]

OpenStudy (studygurl14):

b is the other answer

OpenStudy (immanuelv):

b is one of them... not d i messed up the :D sorry :(

OpenStudy (bossimbacon):

so a ahhhh

OpenStudy (bossimbacon):

tyyy guys

OpenStudy (immanuelv):

^_^ now im gonna be in trouble

OpenStudy (studygurl14):

because.... 2(7x - 10) 14x - 20

OpenStudy (studygurl14):

and \[14x - 18 \neq 14x - 20\]

OpenStudy (anonymous):

6(4x – 3) – 10x = 24x-18-10x= 14x-18 now we can try the choices to knew which are true A. 14x – 15 14x-18=14x-15 14x-14x=-15+18 which gives that 0=3 which is impossible so we have no solution so A is one of the answers B. 2(7x – 10) 14x-18=14x-20 14x-14x=-20+18 which gives that 0=-2 which is impossible so we have no solution so B is the second answers so the answers are A and B to make sure that the answers are A and B let us try to see C and D C. 15x – 14 14x-18=15x-14 -18+14=15x-14x -4 =x so we have exactly one solution so C is not one of the answers D. 2(7x – 9 ) 14x-18=14x-18 14x-14x=-18+18 0=0 so we have infinitly many solutions so D is not one of the answers so we are sure that the answers are A and B

OpenStudy (anonymous):

6(4x – 3) – 10x = 24x-18-10x= 14x-18 now we can try the choices to knew which are true A. 14x – 15 14x-18=14x-15 14x-14x=-15+18 which gives that 0=3 which is impossible so we have no solution so A is one of the answers B. 2(7x – 10) 14x-18=14x-20 14x-14x=-20+18 which gives that 0=-2 which is impossible so we have no solution so B is the second answers so the answers are A and B to make sure that the answers are A and B let us try to see C and D C. 15x – 14 14x-18=15x-14 -18+14=15x-14x -4 =x so we have exactly one solution so C is not one of the answers D. 2(7x – 9 ) 14x-18=14x-18 14x-14x=-18+18 0=0 so we have infinitly many solutions so D is not one of the answers so we are sure that the answers are A and B